A rectangle is 5 feet longer than it is wide. The perimiter of the rectangle is 24 feet. What is the length of the rectangle?
Answer by checkley77(12569)

L=W+5
2L+2W=24
2(W+5)+2W=24
2W+10+2W=24
4W=24-10
4W=14
W=14/4
W=3.5 FT. IS THE WIDTH.
L=3.5+5=8.5 IS THE LENGTH.
PROOF:
2*8.5+2*3.5=24
17+7=24
24=24